### v2

# The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals

Abstract.
The voluntary donation of private health information for altruistic purposes, such as supporting research advancements, is a common practice. However, concerns about data misuse and leakage may deter people from donating their information. Privacy Enhancement Technologies (PETs) aim to alleviate these concerns and in turn allow for safe and private data sharing. This study conducts a vignette survey ($N=494$) with participants recruited from Prolific to examine the willingness of US-based people to donate medical data for developing new treatments under four general guarantees offered across PETs: data expiration, anonymization, purpose restriction, and access control. The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and controls for the impact of confounds including demographics and two types of data collectors: for-profit and non-profit institutions.
Our findings reveal that respondents hold such high expectations of privacy from non-profit entities a priori that explicitly outlining privacy protections has little impact on their overall perceptions. In contrast, offering privacy guarantees elevates respondents’ expectations of privacy for for-profit entities, bringing them nearly in line with those for non-profit organizations. Further, while the technical community has suggested audits as a mechanism to increase trust in PET guarantees, we observe limited effect from transparency about such audits. We emphasize the risks associated with these findings and underscore the critical need for future interdisciplinary research efforts to bridge the gap between the technical community’s and end-users’ perceptions regarding the effectiveness of auditing PETs.
Privacy Guarantee, Health Data Donation, Privacy Expectation, Auditing
ccs: Security and privacy Social aspects of security and privacy
1. Introduction
The altruistic use of personal health data holds immense potential for societal benefit. For example, the public health sector can leverage individuals’ medical histories and health data to analyze epidemiological trends, enhance disease surveillance and risk prediction, diagnose rare and emerging diseases, and accelerate the development of new treatments (Bietz et al., 2015; Aitken et al., 2016; Bartlett et al., 2018; Greene et al., 2019; Kim et al., 2015; Trinidad et al., 2020). Beyond these applications, there are significant economic advantages. The European Commission estimates that innovations enabling greater data sharing and availability could save billions of euros within the EU health sector (Commission, [n. d.]). As a result, the demand for personal health data continues to grow (Wilbanks and Friend, 2016; Bietz et al., 2015), particularly in light of the recent surge in large language models (LLMs), whose effectiveness heavily depends on the data they are trained on.
Upon the acceptance of data-sharing agreements, personal health data is routinely collected by healthcare providers (Bartlett et al., 2018; Kim et al., 2015) as well as personal mobile apps and devices (Greene et al., 2019). This data is often accessed by third parties through data-sharing agreements that govern data donation practices (Trinidad et al., 2020). However, people frequently self-report that privacy concerns deter them from donating their sensitive information (Redmiles, 2020; Voigt et al., 2020; Silber et al., 2022; Greene et al., 2019; Kim et al., 2015). For instance, many individuals express discomfort with hospital policies that permit data-sharing agreements with third parties without obtaining explicit consent from the data owners (Lewis, 2019).
In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage (Popa et al., 2011; Pappas et al., 2014), data processing (Arasu et al., 2013; Bajaj and Sion, 2011; Baumann et al., 2015; Gupta et al., 2019; Chowdhury et al., 2019; Al-Rubaie and Chang, 2019; Gaballah et al., 2024), and machine learning (ML) training (Wu et al., [n. d.]; De Viti et al., 2022). For example, CoVault (De Viti et al., 2022) enables multiparty analytics for epidemics-related datasets while preserving data confidentiality, and Anonify (Gaballah et al., 2024) supports decentralized, dual-level anonymity for secure medical data donation.
However, in order for such PETs to impact the availability of data, the guarantees they offer must impact user privacy concerns significantly enough to alter their willingness to donate their data.
In order to evaluate whether the guarantees offered by PETs ultimately impact people’s willingness to donate their data, at least three steps are necessary. First, we must be able to effectively explain guarantees (either general guarantees or those of a particular PET) to the people they aim to protect: end-users. Prior work has focused on doing so for particular PETs, chiefly differential privacy (DP) and end-to-end encryption (E2EE) (Cummings et al., 2021; Abu-Salma et al., 2018; Akgul et al., 2021; Bai, 2019; De Luca et al., 2016; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023). Second, using these explanations, we must evaluate whether the guarantees impact people’s privacy concerns (or expectations), and in turn, their willingness to share data. Prior work has done such evaluations using self-report studies focused on individual PETs, again primarily DP (Cummings et al., 2021; Bullek et al., 2017; Nanayakkara et al., 2023) or E2EE (Dechand et al., 2019; Bai et al., 2020; Abu-Salma et al., 2018)). Third, while rare, we must validate those results in the field to confirm that the effects observed replicate when studying actual vs. intended behavior (e.g., (Dooley et al., 2022b; Acquisti et al., 2013)).
Instead of focusing on a singular PET and it’s accompanying guarantees, we focus on addressing steps one and two for the general security and privacy guarantees that recur across PETs:
  1. PG(1):
    Anonymization (Murthy et al., 2019; Shahid et al., 2022): data is not linkable to its owner (as defined in (Chevrier et al., 2019));
  2. PG(2):
    Access control (Greene et al., 2019; Hu et al., 2006; Ni et al., 2010; Tourani et al., 2017): data is accessible only by authorized people, specified in the data collection agreement;
  3. PG(3):
    Data expiration (Politou et al., 2018; Sarkar et al., 2018; Perlman, 2005): data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
  4. PG(4):
    Purpose restriction (De Viti et al., 2022; Byun et al., 2005): data is only used for the stated collection purpose, specified in the data collection agreement;
While previous research has explored how the combined presence of guarantees PG(1)-PG(4) affects data-sharing intentions (Aitken et al., 2016; Mori, 2016), our study takes a different approach by focusing on the impact of each guarantee individually. Our objective is to identify which specific protections or factors most significantly influence privacy expectations and, in turn, reported data-sharing intentions.
Additionally, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
  1. AG(1):
    Expert auditing (Li et al., 2015; Wang et al., 2011): engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
  2. AG(2):
    Self auditing (Holt and Seamons, 2005; Chen et al., 2022): granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
We examine the role of these PET guarantees and their audits in shaping privacy expectations and influencing data-sharing intentions within a specific context used in prior work (see e.g., (Cummings et al., 2021; Nanayakkara et al., 2023)): medical data donation. We ask:
  1. RQ(1):
    How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?
  2. RQ(2):
    How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?
To address these questions, we conduct a vignette survey ($n=494$) following the best practice methodology established by prior work (Cummings et al., 2021; Abu-Salma et al., 2018; Xiong et al., 2020). Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with information on how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)). Furthermore, we isolate the effect of PG(1)–PG(4) and AG(1)–AG(2) from external confounding factors identified by prior work:
  • The data-collection entity (Trinidad et al., 2020; Hendricks-Sturrup and Lu, 2023): whether the collecting entity is a for-profit or a non-profit organization;
  • Socio-demographics (Hargittai et al., 2020; Dooley et al., 2022a; Trinidad et al., 2020): age, gender, education, donation history, and technical background.
We find that even when told nothing about PETs implemented by the entity, participants are 23% more likely, on average, to expect a non-profit to implement PG(1)–PG(4) and AG(1)–AG(2). As a result of these already high privacy expectations for non-profit organizations, we find that mentioning a specific privacy protection in the survey does not significantly enhance people’s willingness to donate towards non-profit entities: even when no privacy protection is explicitly mentioned, 89% of the participants are willing to donate to a non-profit entity. In contrast, for-profit entities need to effectively demonstrate their privacy protections; indeed, explicitly mentioning privacy protections in the survey does increase privacy expectations of for-profit entities from 50% to the level of non-profit entities. Privacy expectations, in turn, influence the willingness to donate.
Furthermore, while the technical community has suggested external audits as a mechanism to increase trust in PET implementation, our initial inquiry suggests that more work is needed to align the perceptions of the experts and end users about the effectiveness of such audits. In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
We argue that it is critical for non-profit entities to rigorously implement data privacy measures, as any future data leak could lead to a significant loss of trust. Prior research (Acquisti et al., 2013) underscores that users place greater value on maintaining their expected privacy than on gaining additional privacy they did not initially anticipate. In contrast, we highlight the risk of for-profit entities engaging in ”privacy washing” (Cirucci, 2024), where statements about PETs are used to artificially raise privacy expectations to encourage data collection. At the same time, our findings reveal that respondents are perceptive of the limitations in general PET guarantees, particularly concerning protections against data breaches. This underscores the need for future research to explore how to effectively communicate the potential of stronger, emerging PET guarantees and address the skepticism of end-users.
2. Related Work
In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of privacy-enhancing technologies (PETs), and educational or explanatory strategies related to PETs.
[…]
Data donation.Data donationData donation. The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments. Indeed, the COVID-19 pandemic has further intensified the demand for personal health data (Scott et al., 2020). Despite the importance of this data, its sensitive nature poses significant obstacles to its collection (Fahey and Hino, 2020). Prior work indicates that individuals are generally willing to donate their data for altruistic purposes (Pilgrim and Bohnet-Joschko, 2022; Aitken et al., 2016; Seberger and Patil, 2021), though there is a notable reluctance when it comes to their health data specifically (Middleton et al., 2020; Mori, 2016). However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation (Goodman et al., 2017; Bartlett et al., 2018).
[…]
Privacy concerns.Privacy concernsPrivacy concerns. Concerns about the privacy and the misuse of donated data are prominent among potential donors (Mori, 2016). These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked (Weitzman et al., 2012; Aitken et al., 2016; Kim et al., 2015; McCormick and Hopkins, 2021). A common source of these concerns is distrust of the receiving entity (Weitzman et al., 2012; Howe et al., 2018; Aitken et al., 2016; Stockdale et al., 2019), often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties (Kacsmar et al., 2022; Middleton et al., 2020). This distrust is intensified when participants are unfamiliar with the recipient entity (Weitzman et al., 2012; Howe et al., 2018), especially if it is a governmental or for-profit organization (Aitken et al., 2016; Stockdale et al., 2019; Garrison et al., 2016; Hendricks-Sturrup and Lu, 2023; Khanra, 2023; American Medical Association, 2022). Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling (Seh et al., 2020). People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information (Seh et al., 2020).
[…]
Mental models on PETs.Mental models on PETsMental models on PETs. PETs aim to address user privacy concerns. While various PETs are available for health data donation environments (Jordan et al., 2022), prior work indicates that participants with greater online privacy literacy tend to have more trusting beliefs in PETs (Harborth and Pape, 2020). Studies focusing on sensitive personal data, including health data, reveal similar findings. For instance, Dechand et al. (Dechand et al., 2019) found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access. Furthermore, Lerner et al. (Lerner et al., 2022) noted that some participants doubt the existence of true privacy.
Non-experts may misunderstand privacy technologies (Wu and Zappala, 2018; Oates et al., 2018; Dechand et al., 2019), e.g., viewing encryption merely as restrictive access control (Wu and Zappala, 2018) or confusing it with data encoding (Abu-Salma et al., 2017). Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections (Tang et al., 2021; Velykoivanenko et al., 2022; Abu-Salma et al., 2018; Gerber et al., 2019; Pham and Castro, 2019; Vaniea et al., 2014). As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality. We investigate the alignment between stated privacy guarantees and people’s expectations for how their data will be protected as part of RQ1.
[…]
Explaining PETs.Explaining PETsExplaining PETs. To address these misunderstandings and evaluate the impact of PETs on downstream factors such as privacy concerns (or expectations) and willingness to share data, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts (Cranor et al., 2006). Prior work focuses heavily on E2EE and DP (Abu-Salma et al., 2018; Bai, 2019; Cummings et al., 2021; De Luca et al., 2016; Akgul et al., 2021; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023). These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge. Methods found effective to explain PETs include visualizations (Stransky et al., 2021), mental models (Stewart and Lacey, 2012; Lin et al., 2012), nutrition labels (Kelley et al., 2010), metaphors (Karegar et al., 2022), short statements (Papacharissi and Fernback, 2005), and privacy games (Tupsamudre et al., 2018; Karl and Tao, 2023). In this work we use short statements to briefly explain the relevant guarantees in order to evaluate the impact of these guarantees on participants’ privacy expectations and willingness to donate medical data.
[…]
Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing. Prior work on the effect of PETs and their accompanying PGs typically focuses on the effect of a combination of guarantees offered by a specific PET (e.g., DP (Valdez and Ziefle, 2019; Cummings et al., 2021) or E2EE (Dechand et al., 2019)) on privacy concerns . A more limited subset of studies evaluate the impact on willingness to share data (Valdez and Ziefle, 2019).
Instead of focusing on a specific PET and it’s combined accompanying PGs we seek to examine and compare four PGs that are shared across many PETs: data expiration, data anonymization, use restriction, access control, and auditing (see § 1). In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity. Furthermore, we focus on the effect of two different auditing processes – expert auditing and self auditing – on enhancing privacy expectations and willingness to share data. Other studies have examined public understanding of data retention (Murillo et al., 2018) and general data anonymization (Haddow et al., 2011), and one study examines willingness to donate medical data if it is allowed to be used for secondary purposes (Juga et al., 2021). To our knowledge, no prior work focuses on comparing the effects of the four core PGs we study nor the effects of auditing PETs on user privacy expectations or willingness to donate data.
3. Methodology
As mentioned in § 1, we address research questions RQ1 and RQ2 through a user survey. In this section, we discuss the ethical considerations of our study, outline the survey design (§ 3.2), and describe the cognitive interview process used to ensure participants’ understanding of the presented scenarios (§ 3.3). We also describe our participant selection strategy (§ 3.4) and the limitations of our methodology (§ 3.6).
3.1. Ethics
Before we began our research, our ethics review board approved all of the study methods, including the cognitive interviews and the survey. A consent form detailing the data storage, handling, and survey compensation is presented at the beginning of the survey. Participants can opt out of disclosing any demographic data without influencing their compensation in the survey.
Each participant received $1.20\$$ upon completing the survey. To determine this figure, we ran a test of 20 participants, who took 5m56s (median) to complete it. Thus, the actual compensation was $12\$$/hr, which aligns with Prolific recommendations. For the signup survey of the cognitive interview, we set the study name to “Sign Up for Paid Interview on Data Donation” with a payment of $1.25\$$ ($13.68\$$/hr based on a loose completion estimate of 5 minutes). However, most respondents refused our interview request. Thus, in order to encourage participants interested in attending the interview to complete the screener, we lowered the compensation for the screening survey to $9.96\$$/hr. The interview reward is $15\$$, and all interviews lasted at most thirty minutes.
The survey is conducted through Qualtrics (Qualtrics, 2020), with the IP address collection feature disabled to protect respondents’ privacy. Interviews are conducted via Zoom (zoom, 2024), and participants are instructed to avoid using their real names when joining the meeting to ensure anonymity. Research participants are anonymously recruited through Prolific (Prolific, 2024), as described in greater detail below.
3.2. Survey Structure
Figure 1. Survey design. Each participant is presented with a donation scenario and a survey introduction with two variations: recipient entity and disease. statements (PG(1)–PG(4) and AG(1)–AG(2)) are also presented in the donation scenario. Participants are then asked five sets of questions (survey questions). When the control statements are presented, the trust (dashed) is omitted. The willingness to donate and trust questions (dark-shaded) are associated with an open-text question. The scenario understanding is used as a filter. Participants who did not understand the donation scenario were excluded from the analysis. A sample screen that participants would encounter is provided in Figure 6.
Our survey design is shown in Figure 1, and the full survey can be found in Appendix A. First, we present to the respondents a suvery introduction, which we detail in Figure 2.
Survey introduction
Imagine that an entity wants to develop a new treatment for disease. They need medical data from people with and without disease to develop the treatment. They ask you to donate your medical record to help develop the treatment. Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
Figure 2. Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease. The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) (for Medicare and Services, 2024).
Then, we present to the respondents either a control (no statement) or an experimental statement. Experimental statements are composed of either one privacy statement, alone, or one privacy statement and one auditing statement. Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements. We have written these statements to align with PG(1)–PG(4) (see § 1), and we list them in Figure 3. We further distinguish the auditing guarantee into two different auditing statements to align with AG(1)–AG(2), and we list them in Figure 4.
PG(1): Anonymization — personally identifiable data will not be collected.
The privacy-preserving technology removes any personal identifiable information at the time of data collection, so that the data stored by the recipient entity is not linkable to its data owner (i.e., the data is anonymous).
PG(2): Access control — only relevant scientists can read data.
The privacy-preserving technology restricts data access to the authorized scientists (within the recipient entity) which are working on the treatment for the given disease.
PG(3): Data expiration — data will be deleted later.
The privacy-preserving technology discards the donated data, or makes it inaccessible after a given expiration time, which can be chosen by the data donor in the data collection agreement.
PG(4): Purpose restriction — data will not be used for other purposes.
The privacy-preserving technology ensures that the recipient entity can only use the donated data to develop a treatment for the given disease, and not in the context of any other research they may be working on (e.g., different disease).
Baseline: Control
No privacy-preserving guarantee is at all mentioned.
Figure 3. Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
AG(1): Expert auditing — the implementation of PET will be verified by experts.
An external advisory board of scientists and software engineers appointed by the recipient entity will regularly verify that the privacy-preserving technology is working as described. The results of this verification will be made public.
AG(2): Self-auditing — the implementation of PET can be verified by anyone.
Anyone interested, including the respondent and the experts the respondent trusts, will be able to verify that the privacy-preserving technology is working as described. Anyone, including the respondent and the experts the respondent trusts, can make their verification results public.
Baseline: Control
No auditing is at all mentioned.
Figure 4. Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
Our survey questions assess respondents’ self-reported:
  • scenario understanding, used to filter out respondents that report not understanding the scenario (see Section 3.4): “How would you rate your understanding of the above scenario?” (4 point Likert scale: Fully understand - Not understand).
  • willingness to donate their health data to the recipient entity: “In this scenario how likely would you be to donate your medical record?” (4 point Likert scale: Very likely - Very unlikely).
  • privacy expectations regarding the specific privacy guarantees we investigate, measured by their agreement with the statements presented in Figure 5 using the same Likert scale as willingness to donate. [^1] This question included an additional attention check with a sixth statement: “If I donate my data, I will meet Albert Einstein.” Respondents who did not answer ‘Very Unlikely’ were removed from the experiment, as detailed in Section 3.4. We assess participants’ expectations about all guarantees, regardless of what statement they were presented.
  • trust that the recipient entity will implement the protection described by the privacy statement: “I trust the entity will handle my data as described.” (4-point Likert scale Strongly agree - Strongly disagree).
  • demographics and experiences, including age, gender, education level, donation history, technical background (i.e., education background and job field relationships with computer technologies), and egocentricity(i.e., whether respondents or their close relatives have the disease, as defined in (Grgić-Hlača et al., 2020; Plane et al., 2017)).
Anonymization
My full name or other personal identifiable information will be linked to the donated medical record.
Access control
Any employee at the recipient entity will be able to access the donated medical records.
Data expiration
The donated medical record will be deleted at a set point in time.
Purpose restriction
The donated medical records will be used for another purpose without my consent
Expert auditing
A group of independent experts will verify whether the privacy-preserving technology works and publish a report on their findings.
Self-auditing
I will be able to hire someone to verify that my medical record is protected as described.
Figure 5. To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
We also ask respondents to explain their responses to the willingness to donate and trust questions. These explanations are collected through open-ended questions: “Why you are willing (or unwilling) to share your medical record with this entity?” and “Please explain why you do (or do not) trust that the entity will handle your data as described.” We note that the scenario understanding and trust questions are only asked when a non-control privacy statement is presented.
A sample screen that participants would encounter is provided in Figure 6.
Imagine that a non-profit organization (e.g., university, non-profit research institute) wants to develop a new treatment for Hypertension (High blood pressure). They need medical data from people with and without Hypertension (High blood pressure) to develop the treatment. They ask you to donate your medical record to help develop the treatment. Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
The non-profit organization uses a privacy-preserving technology to ensure that your data is deleted after a set period of time. You will be able to choose from a list of time period options how long your data will be stored. Anyone interested, including you and experts you trust, will be able to verify that the privacy-preserving technology is working as described. Anyone, including you and experts you trust, can post their verification results publically.
How would you rate your understanding of the above scenario? Fully understand Mostly understand Partially understand Not understand
Figure 6. Sample survey screen that composed of three segments. The first segment presents the survey introduction, which varies the disease and the entity among all participants. The second segment includes the randomized privacy statement and auditing statement, with color coding applied to emphasize these elements. The third segment contains the survey question that participants are required to answer.
3.3. Survey Refinement
The final survey was refined through multiple iterations. Before launching the survey, we prescreened 49 users on the Prolific online platform (Prolific, 2024) by asking them to complete a short demographics survey and indicate their willingness to participate in a cognitive interview. Then, we conducted 17 cognitive interviews to ensure participants understood the survey correctly. Each interview involved presenting the survey to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements. We additionally presented all statements to the interviewees and asked for their understanding and perception of each of them. We continued refining the survey through cognitive interviews until no further constructive feedback was received. In the final interview, we observe a clear understanding of the statements and the donation scenario. For example, the participant described anonymization (PG(1)) as ”eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked” and access control (PG(2)) as ”some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in.” Our final survey can be found in Appendix A.
3.4. Survey Sampling
Following the recommendation of (Tang et al., 2022), we recruited participants through Prolific, gathering responses from 560 individuals. To ensure focused results, we restricted respondents to adults residing in the U.S. and requested a gender-balanced distribution. Our recruitment included 272 men, 275 women, 9 nonbinary individuals, and 4 participants who chose not to disclose their gender.
Of these 560 respondents, we sequentially excluded the following from the analysis: 35 that submitted incomplete responses; 20 that failed our attention check (§ 3.2); and 11 that indicated they did not understand the survey scenario (using scenario understand). Thus, our final dataset comprises 494 respondents. To enhance statistical power, we limit the levels of age and education to binary groups based on the data distribution among our participants. We report their demographics in Table 1 and the number of respondents assigned to each condition in Table 2.
Description Category $n$
Age 18 - 29 120 24.3%
30 - 49 222 44.9%
50 - 64 111 22.5%
65+ 41 8.3%
Gender Woman 254 51.42%
Man 231 46.76%
Others 9 1.82%
Education B.S. or above 358 72.47%
Up to H.S. 136 27.53%
Technical Background Yes 129 26.11%
No 365 73.89%
Donation History Yes 56 11.34%
No 438 88.66%
Egocentricity Yes 201 40.69%
No 293 59.31%
Table 1. Participant demographics. We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis, except for “age”, which is treated numerically.
For-Profit Non-Profit
AG(1) AG(2) Ctrl. AG(1) AG(2) Ctrl.
PG(1) 22 18 19 19 20 18
PG(2) 17 21 19 18 18 19
PG(3) 18 20 18 19 19 17
PG(4) 21 18 19 20 18 19
Ctrl. - - 19 - - 21
Table 2. Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 7).
3.5. Analysis
We analyzed the open-text questions about willingness to donate and trust (see § 3.2) using inductive-thematic open coding. Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.2%) responses. Then, they composed a final codebook and double-coded all responses. Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) (McDonald et al., 2019).
In addition to presenting a descriptive analysis detailing the distribution of responses on the survey items that address our research questions, we construct logistic regression models to analyze factors related to two dependent variables: privacy expectations (RQ1) and willingness to donate (RQ2). For RQ1, the independent variables were the presence of a given privacy and/or auditing statement. The dependent variable was the privacy expectation corresponding to the privacy statement presented to a given respondent. Using this model we compare the privacy expectations of respondents in the experimental conditions (those shown a privacy statement) with those of the control groups. Responses from respondents in an experimental conditions are only modeled in the analysis of the privacy statement they were shown.
For RQ2, the dependent variable was willingness to donate, and the independent variables were the presence of a given privacy statement, the privacy expectation for each guarantee, as well as demographics and experiences. We collapsed the independent variables into a binary measure, designating responses of “Likely” or “Very Likely” as True and all others as False. We did so to avoid the ambiguity introduced by intermediate scale points and to simplify the statistical model.
We categorized education into two groups: with and without a bachelor’s degree. We took “no bachelor’s degree” and “less than forty years old” as the reference categories, respectively. For gender, we took man as the reference category. Technical background and donation history are binary factors, and we took the negative response as the reference category. We built separate regression models for the two recipient entities: for-profit and non-profit.
3.6. Limitations
Although the four privacy-preserving guarantees examined in this study (§ 1) cover a broad range of PETs, they are not fully comprehensive and do not capture the full complexity of PETs or real-world threats, which may involve privacy issues beyond the scope of our investigation. The randomly presented disease may also influence the results. Furthermore, our approach may not fully reflect the intricacies of real-world donation scenarios, where multiple guarantees often coexist. While such complexities lie outside the primary focus of our research, qualitative investigations have been conducted in prior work (Aitken et al., 2016; Mori, 2016), and quantitative exploration of these aspects remains an interesting direction for future research.
Despite our efforts to mitigate misunderstandings—through rigorous cognitive interviews, filtering respondents who reported not fully understanding the scenarios, and controlling for privacy expectations in our statistical analyses—the brief descriptions of each guarantee might still have led to partial or incorrect understanding among respondents. This limitation may have influenced our results.
Additionally, while our study employs vignette scenarios that effectively mirror real-world behaviors (Redmiles et al., 2017), there may still be discrepancies between participants’ expressed privacy concerns and their actual behaviors, a phenomenon commonly referred to as the privacy paradox (Barth and de Jong, 2017).
Furthermore, statistically significant results do not inherently eliminate the risk of underlying biases that could skew the findings. Various unmeasured factors not accounted for in our analysis might have influenced the interpretations of the results. Lastly, our respondents were recruited via Prolific, which may limit the generalizability of our findings. The sample might not fully represent the diversity of the U.S. population, nor does it capture the perspectives of individuals from other countries, highlighting a key limitation of this work.
4. Results
4.1. RQ1: Understanding of PGs
Next, we answer the first research question (RQ1, see § 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?”. We analyze and report the survey responses and evaluate the rationales for the responses.
4.1.1. Analysis Setup
We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 3) and auditing statements (listed in Figure 4). In Figure 7, we show the privacy expectations of respondents who were shown different privacy statements. To assess the statistical significance of the descriptive quantitative results presented above, we used binomial logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations. We summarize results, which we distinguish for non-profit and for-profit entities, in Figure 8.
Figure 7. Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario. The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented. The 3rd column (“overall”) reports results across both entities. For example, the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected their data to be anonymized. The agreements are binarized here to present the overall tendency.
Figure 8. Odd ratios and 95% confidence intervals of presenting different privacy statements on the respondents’ privacy expectations toward the two kinds of recipient entities, for-profit and non-profit.
4.1.2. Response Distribution and Statistical Analysis Results
We find that privacy expectations differ based on the data-collecting entity. In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, use restriction to a specific purpose, expert auditing, and self-auditing. In contrast, a higher percentage of respondents (43%-76%) expected non-profit entities to employ these mechanisms, except for self-auditing (AG2). For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): ”I also assume that the data is looked at in the aggregate and likely no one at the company knows me”. In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: ”I trust that the organization will uphold strict privacy and ethical standards”.
In the for-profit scenario, all privacy statements, except for the access-control statement (PG2), significantly increase the corresponding privacy expectations ($p$-value ¡ $0.05$). In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration (PG3) significantly increases expectations (from 57% to 100%). Comparing the for-profit and non-profit subsets of the descriptive results (1st and 2nd column of Figure 7), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios. Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (PG1, 68% and 67%) and access control (PG2, 58% and 42%). It also reduced the gap in expectations for data expiration (PG3, 67% and 100%)) and purpose restriction (PG4, 74% and 84%)).
In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value ¡ $0.05$). However, we note that incorporating an audit statement does not markedly change privacy expectations. The statistical significance of the auditing statements in Figure 8 indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection. As observed in  Figure 7, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
4.1.3. Qualitative Analysis of Privacy Expectations via Open-Answer Responses
Across for-profit and non-profit scenarios, 69% and 85% of respondents, on average, expressed that their positive privacy expectations were formed based on the presence of the privacy statement, belief in privacy obligations of the entity, trust in the entity, or trust in the auditing process.
[…]
Satisfaction with the privacy guarantees..Satisfaction with the privacy guarantees.Satisfaction with the privacy guarantees.. 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement. Some respondents, like P56, vaguely stated that the policy looks promising: ”I see their policy, and they have to follow their own policy”. Others felt explicitly safer and trusted the statement. For example, P494 noted: ”Trust is very important when it comes to medical data. I believe the organization has privacy policies that outline that they collect and will use my data. I also believe the organization will employ security measures to safeguard data”. Similarly, P555 highlighted the presence of the privacy statement: ”(…) purposely states there is software in place to conserve privacy”. and P516 felt their information was safe because of it: ”(…) with the privacy protection in place, they are isolating the data they need while basically ”throwing out” the rest by putting in under that protection. In essence, my information is safe, and they’re only using what they said they’d use”. P208 also felt safer contributing to research knowing the privacy mechanisms were in place: ”I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above”.
[…]
Belief in legal and reputational obligations..Belief in legal and reputational obligations.Belief in legal and reputational obligations.. 16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements and reputational concerns. For example, P28 stated regarding for-profits: ”A for-profit organization wouldn’t want to violate HIPAA, HITECH laws”. Additionally, P117 wrote ”I would trust them to do the right thing so they won’t face lawsuits”. Regarding non-profits, P141 noted: ”I trust that the law will restrict any data leaks to third parties”.
[…]
Trust in the recipient entity..Trust in the recipient entity.Trust in the recipient entity.. 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons. For instance, P70 succinctly said ”I feel they are reliable and trustworthy” and P50 stated that ”I assume they take their research seriously, so they would handle the data carefully”. Other respondents, like P145, reported having no reason not to trust it: ”I have no reason to think they would do anything nefarious with my medical data”.
[…]
Trust in the auditing process..Trust in the auditing process.Trust in the auditing process.. 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability. For example, P202 showed some reservation but found confidence in the public nature of audits: ”I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public. That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute”. Additionally, P540 had trust in the entity’s data handling because ”they let outsiders audit them” and P51 (who received the expert-auditing statement) because of ”safelocks and checks in place”. On the other hand, one interviewee doubted the expert auditing process and stated: ”somebody else says something doesn’t mean that it’s real”. No survey respondent explicitly raised concern towards the auditing process.
Across for-profit and non-profit scenarios, 31% and 15% of respondents, on average, expressed negative privacy expectations because they were skeptical of the privacy statement or doubted whether the recipient entity would actually employ PETs as stated. Their qualitative responses offer insights on the underlying reasons:
[…]
Skepticism on the privacy statement and limits of privacy-preserving guarantees..Skepticism on the privacy statement and limits of privacy-preserving guarantees.Skepticism on the privacy statement and limits of privacy-preserving guarantees.. 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement. In the non-profit scenario, P34 stated that ”no privacy technology is foolproof”. Similarly, in the for-profit scenario, P45 wrote: ”I don’t trust that the privacy-preserving tech would work.” More precisely, privacy statements were considered too ideal to be fully enforced. Some respondents believed that unauthorized employee access would be unavoidable. For instance, P71 said: ”Why would I trust someone other than my doctor with my medical records? These days especially, I don’t trust anyone. There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records. Insane!”. Some others felt that data breaches and cyberattacks were inevitable. For instance, P83 stated that ”the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked”, while P124 noted that ”corporate data breaches are very common”, and P80 echoed this sentiment, saying: ”I believe the [intentions] will be good, but data can be hacked”. We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario. For example, P474 mentioned: ”They can check their privacy technology all they want but when there is a breach it is done and info is stolen. After it fails then they say sorry and offer monitoring but the info is still stolen”.
[…]
Doubt on the recipient entity’s motivation to employ PETs..Doubt on the recipient entity’s motivation to employ PETs.Doubt on the recipient entity’s motivation to employ PETs.. 8.6% and 3.3% of the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all. Most criticism was directed at for-profit entities. For instance, P55 noted that for-profit entities ”will do what is profitable and not much more than that”. Similarly, P119 stated ”(…) because it is a for-profit organization. I expect them to cut corners” and P32 wrote that ”for-profit organization have low standard of morality”. However, some respondents also expressed concerns about non-profit entities. For example, P190 in the non-profit scenario remarked: ”Medicine has become a business. My data is only useful to them if it helps them make more money. Money comes first before the actual well being of humans”.
4.2. RQ2: Willingness to donate health data
Next, we answer the second research question (RQ2, see § 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?”. As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
4.2.1. Analysis Set Up
In Figure 9, we summarize the respondents’ willingness to donate their health data to a recipient entity. The 1st column (“overall”) reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity. The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
We constructed a logistic regression model to understand the factors that influence willingness to donate. We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models. The results are summarized in Table 3.
In Figure 10, we visualize the overlap in experimental group respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described.
4.2.2. Response Distribution and Statistical Analysis Results
Figure 9. Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario (e.g., the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.) The agreements are binarized here to present the overall tendency.
Figure 10. Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded). The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (310). $N=454$.
Entity Factor Levels OR 95% CI $p$-value
Both Entity - FP 0.38 [0.24, 0.59] <0.001
For-Profit Privacy Statement
   PG(1) 0.65 [0.25, 1.70] 0.381
   PG(2) 0.68 [0.27, 1.75] 0.429
   PG(3) 0.86 [0.32, 2.28] 0.756
   PG(4) 0.62 [0.22, 1.70] 0.352
Privacy Expectation
   PG(1) 1.09 [0.66, 1.80] 0.741
   PG(2) 1.60 [0.96, 2.69] 0.073
   PG(3) 1.70 [1.01, 2.88] 0.047
   PG(4) 3.25 [1.78, 5.87] <0.001
   AG(1) 2.46 [1.40, 4.27] 0.002
   AG(2) 1.32 [0.76, 2.27] 0.317
Demographics & Experiences
   Education 0.73 [0.43, 1.26] 0.264
   Age 1.00 [0.98, 1.01] 0.765
   Gender 1.30 [0.80, 2.12] 0.304
   Tech Background 1.57 [0.88, 2.80] 0.127
   Egocentricity 0.69 [0.42, 1.14] 0.144
   Donation History 3.25 [1.31, 8.01] 0.011
Non-Profit Privacy Statement
   PG(1) 1.27 [0.52, 3.13] 0.603
   PG(2) 0.94 [0.37, 2.38] 0.901
   PG(3) 1.17 [0.45, 3.00] 0.746
   PG(4) 1.54 [0.61, 3.90] 0.369
Privacy Expectation
   PG(1) 1.72 [1.02, 2.89] 0.043
   PG(2) 1.46 [0.87, 2.46] 0.152
   PG(3) 1.27 [0.73, 2.25] 0.400
   PG(4) 5.21 [2.61, 10.38] <0.001
   AG(1) 2.59 [1.35, 4.90] 0.004
   AG(2) 1.17 [0.68, 2.03] 0.563
Demographics & Experiences
   Education 0.46 [0.26, 0.80] 0.007
   Age 1.00 [0.98, 1.02] 0.897
   Gender 1.16 [0.69, 1.97] 0.560
   Tech Background 0.71 [0.40, 1.26] 0.237
   Egocentricity 1.36 [0.83, 2.23] 0.217
   Donation History 2.25 [1.06, 4.75] 0.033
Table 3. Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
There was no direct relationship between showing a privacy statement and respondents’ willingness to donate. However, several privacy expectations positively correlate with willingness to donate for both non-profit and for-profit entities, except in specific cases like access control for both for- and non-profit and data expiration for the non-profit.Visualized in Figure 10, we observe high alignment among the three explored constructs, where 236 (52.0%) respondents gave all positive responses and 63 (13.9%) respondents gave all negative responses. Combined with the results of RQ(1), the findings suggest that willingness to donate is influenced by privacy expectations, which are, in turn, impacted by privacy statements.
Privacy statements and expectations are not the only reason for donation intentions however. While participants were generally less willing to donate to for-profit entities (OR=0.38, $p=\textless 0.001$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities. This indicates that other non-privacy-related factors, such as the perceived impact or benefits of the donations, also influence participants’ willingness to donate.
Furthermore, we observe that respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.46, $p=0.007$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement. They have heightened privacy concerns that need to be addressed. For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
Finally, we found that prior donors are more willing to donate to both entities. Respondents who have donated before are more willing to donate to both entities, with the OR value of 3.25 ($p=0.011$) and 2.25 ($p=0.033$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments. For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes.” and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection”.
4.2.3. Qualitative Analysis of Donation Intention Via Open-Answer Responses
Many respondents reasons for donating (or not) included a desire to support research, ego-centric connections, or desire for personal reward. In the for-profit scenario, 50.2% of respondents’ reasons for donating (or not) fell into these categories, in the non-profit scenario, 53.3% did.
One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that ”there are a number of things I could share to research such as money, time, data, and more. I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines.” Similarly, P97 stated: ”even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in.”
Ego-centric reasons, particularly familial connections, also motivated donations. P154 mentioned donating because ”my dad died when his heart failed on him, and I’d have rather that not had happened”, and P233 noted that they ”would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did.”
However, their existing (negative) experiences with for-profit healthcare entities also deter some of them from making donations. For example, P136 claimed they have heard a number of times that the entity says they won’t sell your information but can get around it by calling it something else, thus technically selling the information. P244 also mentioned that their information has been leaked before by a clinician.
Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that ”I would give them my data if I got paid for it. I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled.”
An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use. For instance, participants P495, P463, and P459 all share this perception. However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities. For example, P45 ”just lost my father to heart failure thirteen days ago. (…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.
Respondents expressed negative sentiments toward for-profit entities collecting data for research. The desire for personal reward related closely to lack of donation intent for for-profit organizations. In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations. For instance, P191 and P202 respectively stated: ”I have a negative connotation with for profit organization” and ”I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me.” Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further. For instance, P36 wrote: ”For profit organization makes profit using the data. I’m never going to donate it. They can BUY it from me for a reasonable amount.”
Alternatively, a small group of 1.3% respondents explicitly mentioned that they were willing to donate data to entities of non-profit nature, but even they expressed concerns about potential misuse or unethical behavior within these organizations. For instance, P122 mentions: ”it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with. Scares me a bit.” Then, P282 notes: ”it is difficult to police everyone in a non-profit organization. I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have.”
Privacy-related donation considerations focused on data sensitivity and leakage. Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that ”there are so many data leaks from so called safe places. No one can anticipate what hackers can do in the future.” Others cited previous data leak experiences as to why they would not donate, as P290 said ”my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her. For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others.” Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that ”regardless of privacy policy, the probability of a leak or misuse is high. unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.”.
Another reason was that the participant was concerned that their medical data included too much or too specific information. P155 highlighted this concern: ”I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not”; as did P229: ”It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me.” Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., ”I don’t have mayor illnesses or nothing to hide so Im ok with that”) or because they already share information with other entities (e.g., ”Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different”).
5. Concluding Discussion
Like our survey, when soliciting data donations, recipient entities usually provide an introduction motivating their need to collect the individual’s data, outline privacy guarantees, and seek consent for data donation. Privacy statements function similarly to advertising (O’Flaherty, 2022) by shaping and influencing consumers’ expectations of privacy (Schaub et al., 2015). Yet, these statements do not always have impact.
Prior work has highlighted the ways in which such statements can fail due to being vague or inaccurate. Our work highlights how the impact of privacy statements is moderated by people’s pre-existing privacy expectations.
Our participants reported strong pre-existing privacy expectations, even in the absence of explicit privacy statements. We observe that these expectations vary strongly based on the profit model of the entity. For for-profit entities, where baseline privacy expectations are lower, strong privacy statements can effectively elevate these expectations among users. In contrast, the impact on non-profit organizations is minimal because users already hold high privacy expectations for these entities.
[…]
The Halo Effect.The Halo EffectThe Halo Effect. Participants qualitative responses suggest that their strongly positive privacy expectations for non-profit entities are rooted in arguments of morality: non-profits are perceived as more ethical and dedicated to serving the public good because they are not pursuing profit. This findings are consistent with a phenomenon termed “the halo effect”, which describes how an overall positive impression of a particular entity (e.g., the good work that a non-profit does) can lead to ungrounded assumptions about how the organization operates (Nisbett and Wilson, 1977; de Bruin Cardoso et al., 2024). While prior work in the business literature focuses on how the halo effect influences expectations about business practices such as donation management, our work illustrates that the halo effect also influences people’s expectations about the way non-profits handle data privacy.
These positive expectations do not always align with reality. Privacy safeguards are not consistently enforced or widely practiced; for example, many health organizations rely on broad consent frameworks rather than adopting stricter purpose-restricted consent measures (Kaye et al., 2015). This gap between expectations and reality can result in privacy expectation violations, a major source of privacy concerns as explained by the theory of contextual integrity (Nissenbaum, 2009). Indeed, a well-known non-profit mental health helpline shared it’s data with a commercial entity, leading to significant public outcry (Levine, 2022). Future work may seek to explore how to design privacy communications that set appropriate expectations in the presence of such effects, as a prior work has sought to do in other domains (e.g., (Burton et al., 2015)).
[…]
Privacy Washing.Privacy WashingPrivacy Washing. Conversely, our participants’ had far lower initial expectations of the privacy practices of for-profit entities due to their perceived singular focus on profit (a “horn” effect). As a result, information about data protection guarantees significantly raised their privacy expectations, and in turn, their willingness to donate their data.
This means that for-profit entities have the opportunity to gain “competitive advantage” (Schaub et al., 2015) through influencing and enhancing user privacy expectations via transparently implementing and effectively communicating about PETs. However, this effect also posses a vulnerability to ”privacy washing” (MainWP, 2023; Cirucci, 2024), in which organizations make vague or exaggerated claims about PETs to falsely raise privacy expectations. Such deceptive practices can mislead individuals into sharing their data, erode trust, and discourage future data donations if individuals later feel deceived.
[…]
The Role of Auditing.The Role of AuditingThe Role of Auditing. In theory, auditing could offer verification that combats both the halo effect and privacy washing by grounding assumed or stated privacy guarantees in technical reality. Unfortunately, we find that auditing has a limited effect on elevating privacy expectations among participants, except in scenarios involving purpose use restrictions by for-profit entities. Participants understand that while PETs and auditing processes are in place, they are not infallible, and auditing does not substantially mitigate the practical risks associated with data breaches or unauthorized access.
This perception shows a disconnect between the perspectives of experts and participants, as shown in (Dechand et al., 2019; Oates et al., 2018; Wu and Zappala, 2018). Experts often regard auditing as a fundamental assurance mechanism that reinforces trust and compliance. In contrast, participants are more concerned with the overall protection of their data, regardless of specific technical implementations or compliance processes. As one respondent noted, ”They can guarantee privacy all they want; things still get hacked.” True privacy is doubted to exist (Lerner et al., 2022).
[…]
Building Trust.Building TrustBuilding Trust. Our respondent’s prior experiences with data leakage often informed this fatalistic perspective on privacy. Technologies like multi-party computation (MPC), although not yet scalable, have the potential to minimize the impact of a breach by specifically distributing sensitive information across different entities to ensure that no single entity can reconstruct the information on its own. However, deploying such systems requires careful engineering to prevent distributed data breaches on all the entities involved. Existing skepticism about data security means that any failure of these systems could severely damage trust, potentially more so than if simpler, well-understood methods were compromised.
[…]
Future Work.Future WorkFuture Work. In this work, we explicitly examined the influence of four privacy guarantees and two auditing guarantees. Future research could broaden this scope by incorporating a more diverse and comprehensive range of PETs, with a particular focus on understanding how multiple guarantees interact in real-world scenarios where they often coexist. Additionally, researchers can also conduct targeted studies to understand the perspectives of diverse demographic groups, including individuals with a strong distrust of data-handling practices. Identifying and understanding these individuals is critical in fostering confidence in PETs and practices.
Our study focused exclusively on individuals’ perspectives regarding formal medical data donation scenarios. However, medical data are also frequently collected through smart devices and donated with a single click via in-app prompts. Prior work emphasizes the role of modality and timing in the design space of privacy information (Schaub et al., 2015). Future research could explore how such factors intersect with biases such as the halo effect to influence privacy expectations, trust, and willingness to share sensitive information.
Finally, prior work (Seberger and Patil, 2021) studying COVID-19 contract tracing apps – many created through a partnership between for-profits and non-profits – found that people assumed these apps would “operate according to the economics of surveillance capitalism…[making it] unreasonable to expect them to trust the system that profits from the sale of their data.” Future work may seek to further interrogate how such partnerships, which are common in the medical domain, further complicate our understanding of people’s privacy expectations and their sensitivity to statements of data protection.
Appendix A Survey
In this section, we report the full text of our questionnaire. We first ask for a participant themselve or their close relatives have been diagnosed with a disease. In the rest of the questionnaire, we will choose at random one disease the participant has selected, if any, or one disease the participant has not selected. This is labeled in Italic as disease. The entity, also in Italic, may be either non-profit entity or for-profit entity, chosen at random.
Egocentricity
  1. (1)
    Have you ever been diagnosed with any of these diseases?
    • Cancer (Breast, Colorectal, Lung, and Prostate)
    • Diabetes
    • Heart Failure
    • Hypertension (High blood pressure)
    • Stroke
    • None of the above
  2. (2)
    Have your partner and/or close relatives (parents, children, siblings) ever been diagnosed with any of these diseases?
    • Cancer (Breast, Colorectal, Lung, and Prostate)
    • Diabetes
    • Heart Failure
    • Hypertension (High blood pressure)
    • Stroke
    • None of the above
Survey Introduction
Next, we present you with a scenario and ask you how you would respond. Please do your best to imagine yourself in this scenario. There is no right or wrong answer to these questions: we are interested to hear how you would feel about being in this scenario and what you would do.
Imagine that an entity entity wants to develop a new treatment for disease. They need medical data from people with and without disease to develop the treatment. They ask you to donate your medical record to help develop the treatment.
Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
Questions
We will now ask you a series of questions about this scenario. For your reference, the scenario text will appear at the top of each page.
  1. (1)
    How would you rate your understanding of the above scenario?
    • Fully understand
    • Mostly understand
    • Partially understand
    • Not understand
  2. (2)
    In this scenario how likely would you be to donate your medical record?
    • Very likely
    • Likely
    • Unlikely
    • Very unlikely
  3. (3)
    Why your are willing (or unwilling) to share your medical record with this entity?
  4. (4)
    Please rate your agreement with the following statement:
    I trust the entity will handle my data as described.
    • Strongly agree
    • Somewhat agree
    • Somewhat disagree
    • Strongly disagree
  5. (5)
    Please explain why you do (or do not) trust that the entity will handle your data as described.
  6. (6)
    In this scenario, how likely do you think it is that each of the following will occur?
    • My full name or other personal identifiable information will be linked to the donated medical record.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • The donated medical record will be deleted at a set point in time.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • Any employee at the recipient entity will be able to access the donated medical records.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • The donated medical records will be used for another purpose without my consent.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • A group of independent experts will verify whether the privacy-preserving technology works and publish a report on their findings.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • I will be able to hire someone to verify that my medical record is protected as described.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • If I donate my data, I will meet Albert Einstein.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
  7. (7)
    History In real life, have you ever donated your medical record?
    • Yes, I have donated my medical record for disease
    • Yes, I have donated my medical record for research of other illnesses.
    • Yes, I have donated my medical record for non-medical purposes.
    • No, I was asked to donate my medical record but I declined to participate.
    • No, I have never been asked to donate my medical record.
    • I prefer not to answer.
Internet Skills
  1. (1)
    How familiar are you with the following terminology? Please choose a number between 1 and 5 where 1 represents “no understanding” and 5 represents “full understanding” of the term.
    • Advanced search
    • PDF
    • Spyware
    • Wiki
    • Cache
    • Phishing
  2. (2)
    Which of the following best describes your educational background or job field?
    • I have an education in, or work in, the field of computer science, computer engineering or IT
    • I DO NOT have an education in, nor do I work in, the field of computer science, computer engineering or IT.
    • Prefer not to answer.
  3. (3)
    Have you ever heard any of the following techniques? (select all that apply)
    • Differential Privacy
    • End-to-end encryption
    • Public-key cryptography
    • Trusted execution environment (TEE)
    • Pseudo-party computation
    • None of the above
Demographics
  1. (1)
    What year were you born in? (4 digits)
  2. (2)
    To which gender do you most identify? (Select all that may apply)
    • Woman
    • Man
    • Non-binary
    • Prefer not to answer
    • Prefer to self-describe
  3. (3)
    What is the highest level of education you have completed?
    • Less than high school degree (1)
    • High school graduate (high school diploma or equivalent including GED) (2)
    • Some college but no degree (3)
    • Associate’s degree (4)
    • Bachelor’s degree (5)
    • Advanced degree (e.g., Master’s, doctorate) (6)
    • Prefer not to answer (7)
Appendix B Code Book
[…]
Data Donation.Data DonationData Donation.
  1. (1)
    Irrelevant response
  2. (2)
    Opt to donate regardless
  3. (3)
    Opt not to donate regardless
  4. (4)
    Require compensation
  5. (5)
    Only donate to a non-profit entity
  6. (6)
    No donation to a for-profit entity
  7. (7)
    Need more information
  8. (8)
    The provided protection is persuasive
  9. (9)
    Generally, the donation scenario looks good
  10. (10)
    Not satisfied with the provided protection
[…]
Trustworthy.TrustworthyTrustworthy.
  1. (1)
    Trust without a specific reason
  2. (2)
    Irrelevant response
  3. (3)
    Perfect protection does not exist
  4. (4)
    The recipient entity has no reason not to protect the data
  5. (5)
    The recipient entity has no reason to protect the data
  6. (6)
    A non-profit entity is trustworthy
  7. (7)
    A for-profit entity is not trustworthy
  8. (8)
    Need more information
  9. (9)
    The provided protection is persuasive
  10. (10)
    Generally, the donation scenario looks good
  11. (11)
    Not satisfied with the provided statement